OpenStudy (swissgirl):

Prove that the relation is an equivalence relation. The Relation Ton R X R given by (x,y)T(a,b) iff \( x^2 +y^2=a^2+b^2 \). Sketch the equivalence classes of (1,2); of (4,0)

OpenStudy (anonymous):

Copied from wikipedia: A given binary relation ~ on a set A is said to be an equivalence relation if and only if it is reflexive, symmetric and transitive. Equivalently, for all a, b and c in A: a ~ a. (Reflexivity) if a ~ b then b ~ a. (Symmetry) if a ~ b and b ~ c then a ~ c. (Transitivity)

OpenStudy (anonymous):

So, on the set RxR, we need to show reflexivity, symmetry, and transitivity.

OpenStudy (swissgirl):

so its reflexive and syymetric

OpenStudy (swissgirl):

ohhhh and transitivity

OpenStudy (anonymous):

Yeah, I mean, that's obvious. You could write a proof for it pretty easily, I'm guessing.

OpenStudy (swissgirl):

ya i am having problems with teh second part i am not sure what they want

OpenStudy (anonymous):

The second part? Sketch the equivalence classes of...?

OpenStudy (swissgirl):

ya like wth does it want me exactly to do?

OpenStudy (anonymous):

An equivalence relations takes a set and breaks it up into partitions, labeling big chunks of the set as "equivalent." So when they ask about the equivalence class, they want to know what things in the set are being grouped together. In this case, the set is RxR, so ordered pairs. And the way we call things equivalent is if a^2+b^2 is the same. So RxR is getting broken up into equivalence classes based on what the result is when you do a^2+b^2.

OpenStudy (anonymous):

So let's look at the ordered pair (1,2) 

OpenStudy (phi):

I get the impression the equivalence class of (1,2) is the circle on the x,y plane centered at 0,0 with radius sqrt(5)

OpenStudy (anonymous):

That's correct... answer guy.

OpenStudy (swissgirl):

its ok if phi gives me an answer lol like I am studying online so I am teaching all this to myself so sometimes i need to see an example to clarify things

OpenStudy (anonymous):

I'm just fundamentally opposed to giving answers, particularly when someone else is working with someone and then you just hop in and drop an answer.

OpenStudy (swissgirl):

ya thats true

OpenStudy (swissgirl):

wait but i still dont get y that is the answer let me work that thru

OpenStudy (swissgirl):

ohhhh ok i get it

OpenStudy (anonymous):

But anyways, here's the explanation What are we doing when we do a^2+b^2? Well, if we square rooted that, we'd have the distance from the origin. So we're finding the square of the distance.

OpenStudy (anonymous):

So the equivalence relation that has all a^2+b^2 = 5 means that all of the points have to have the same distance from the origin.

OpenStudy (swissgirl):

ok so the equivalence class of (4,0) is a circle with the origin being (0,0) and a radius of 4 lol ok this one is simple

OpenStudy (anonymous):

Yes. It's pretty simple. Ask another one. =)

OpenStudy (swissgirl):

okkk well no questions that r this style but i found smth similar

OpenStudy (swissgirl):

Let R be the relation on Q defined by \( \frac{p} {q} R \frac{s} {t}\) iff pt=qs. Show that R is an equivalence relation. describe all ordered pairs in the equivalence class \( \frac{2} {3} \)

OpenStudy (swissgirl):

idk this one is a lil bit harder

OpenStudy (anonymous):

Well I would say that the proof part is harder than naming the equivalence class.

OpenStudy (swissgirl):

ok so the ordered pairs wld be multiples of (2/3)

OpenStudy (anonymous):

Actually maybe it's not so bad... \(\Large \frac{p}{t} R \frac{p}{t} \)? Yes, because pt = pt. \(\Large \frac{p}{t} R \frac{r}{s} \rightarrow \frac{r}{s} R \frac{p}{t}\)? Yes, because ps = rt \(\rightarrow\) rt = ps

OpenStudy (swissgirl):

okkkkkk

OpenStudy (anonymous):

And transitivity: \(\Large \frac{p}{t} R \frac{r}{s}, and \frac{r}{s} R \frac{q}{v}\) ps = rt, and rv = qs so p/t = r/s and r/s = q/v, therefore p/t = q/v and pv = qt

OpenStudy (swissgirl):

yaaaaaaaaaaa

OpenStudy (swissgirl):

i solved tran

OpenStudy (swissgirl):

transitivity in a backwards manner but it worked too

OpenStudy (anonymous):

And yeah, I think it should just be multiples of 2/3. \(\Large \frac{2}{3} R \frac{a}{b} \rightarrow 2a = 3b\) So 2/3 = a/b, which just meants that a/b is some multiple of 2/3.

OpenStudy (anonymous):

So that gives you an idea of what this equivalence relation is actually doing. It's calling two fractions equivalent if one is a multiple of the other.

OpenStudy (swissgirl):

okkkkkk thannnkkkkssssssssssssssssssssssssssssssssssssssss

OpenStudy (swissgirl):

That was really clear

OpenStudy (anonymous):

yooooouuuuur weeeeeeeeeeelcoooooooome

OpenStudy (anonymous):

you're*

OpenStudy (anonymous):

Good =D